Two circular plates, each with a radius of 4.22 cm, have equal and opposite charges of magnitude 7.452 μC Calculate the electric field between the two plates. Assume that the separation distance is small in comparison to the diameter of the plates. electric field: 1.506 x10 N/C The plates are slowly pulled apart, doubling the separation distance. Again, assume the separation distance remains sma in comparison to the diameter of the plates. What changes occur with the electric field between the plates? 。The electric field decreases by a factor of 2. O The electric field decreases by a factor of 4. O The electric field stays the same. The electric field increases bv a factor of 2 terms of use contact us help
Answer #1

Solution :-

Radius (r) = 4.22cm = 4.22×10^-2 m

q = 7.452 ×10^-6 c

electric field due to a charged capaciotr is E = Q/eoA

where Q = charge

A = area = pi r^2

E = 7.452 *10^-6/(8.85*10^-12 * 3.14 * 4.22*4.22*10^-4)

E = 1.505 *10^8 N/C